A uniform solid disk, a uniform solid sphere, and
Answer #1

Sphere, disk, hoop

we assume all the objects have the same mass and radious. the MI of the sphere is the least and that of the hoop is the largest, they are more inert in that order, the least inert will have more speed or easy to change its state of rest or of motion.

golf ball, bowling ball and marble.

It is 3-way tie as the MI depends on the actual mass and the size or the radious of the sphere.

the least MI object will be the fastest.

speeds :

MI if sphere Is =2MR^2/5

Total KE (roatational + transitional)

KE = Is\omega2/2 +MV2/2 = (2MR^2/5)*(V/R)^2 /2 + MV^2/2 ( \omega = V/R)

       =7MV^2/10

       = Mgh  --- change in PE

Vs = sqrt(10gh/7)

For disk

MI   Id = MR^2/2

KE = Id \omega2/2 +MV2/2 = (MR^2/2)*(V/R)^2 /2 + MV^2/2

= 3MV^2/4= Mgh

Vd = sqrt(4gh/3)

For the hoop

MI    Ih = MR^2

KE = Ih \omega2/2 +MV2/2 = (MR^2)*(V/R)^2 /2 + MV^2/2

      = 3MV^2/2 = Mgh

Vh = sqrt(2gh/3)